Circular No. 8740

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2006 O3–O8, C/2006 P2–P7 (SOHO) Additional Kreutz sungrazing comets have been found on SOHO web-

site images (cf. IAUC 8739; AK = A. Kubczak, SY = S. Yuan). C/2006 O3–O6 were all very small and stellar in appearance, with no tails; C/2006 O3 peaked at mag ∼ 6.5–7, C/2006 O4 and O5 peaked at mag ∼ 7, and C/2006 O6 peaked at mag ∼ 5.5–6. C/2006 O7 peaked at mag ∼ 3.5–4 with a stubby tail in C3 images; in C2 images, it showed a a diffuse head and a very thin tail that peaked in length at ∼ 29′. C/2006 O8 showed a hint of elongation and peaked at mag ∼ 5.5. C/2006 P2 was tiny, faint, and stellar in appearance, peaking at mag ∼ 7–7.5. C/2006 P3 appeared stellar with a hint of elongation, peaking at mag ≈ 6–6.5. C/2006 P4 was teardrop-shaped in C3 images, peaking at mag ∼ 5.5; in C2 images, it was very faint and diffuse with a very slight elongation but no tail. C/2006 P5, C/2006 P6, and C/2006 P7 were all was quite faint, tiny, and stellar in appearance, with C/2006 P7 reaching mag ∼ 7.5.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 O3 July 17.113 7^{h}31.^{m}2 +19^{o}28′ C3  TH 2006-P28
C/2006 O4  20.179 7 41.9 +18 25 C3  HS 2006-P28
C/2006 O5  21.488 7 47.8 +17 47 C3  HS 2006-P28
C/2006 O6  23.571 7 51.7 +17 20 C3  HS 2006-P28
C/2006 O7  25.679 7 55.2 +16 13 C3/2 HS 2006-P32
C/2006 O8 26.488 8 02.6 +17 07 C3 HS 2006-P32
C/2006 P2 Aug. 1.696 8 30.7 +16 35 C3  AK 2006-P32
C/2006 P3 2.138 8 27.9 +15 40 C3 AK 2006-P32
C/2006 P4  2.431 8 24.8 +15 13 C3/2 SY 2006-P33
C/2006 P5  3.179 8 36.5 +16 05 C3  HS 2006-P33
C/2006 P6  3.363 8 35.5 +15 58 C3  HS 2006-P33
C/2006 P7  4.196 8 39.5 +15 51 C3  AK 2006-P33

COMET C/2006 M1 (LINEAR)

Improved orbital elements from MPEC 2006-P16:

Epoch = 2007 Mar. 1.0 TT

T = 2007 Feb. 13.9184 TT  ω = 122.^{o}8921
e = 0.976913  Ω = 231.6153 }2000.0
q = 3.556288 AU  i = 54.8770
2006 August 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green